# Enhancing Large Vision Language Models with Self-Training on Image Comprehension

### v1

@@ Abstract
Large vision language models (LVLMs) integrate large language models (LLMs) with pre-trained vision encoders, thereby activating the perception capability of the model to understand image inputs for different queries and conduct subsequent reasoning. Improving this capability requires high-quality vision-language data, which is costly and labor-intensive to acquire. Self-training approaches have been effective in single-modal settings to alleviate the need for labeled data by leveraging model’s own generation. However, effective self-training remains a challenge regarding the unique visual perception and reasoning capability of LVLMs. To address this, we introduce Self-Training on Image Comprehension (STIC), which emphasizes a self-training approach specifically for image comprehension. First, the model self-constructs a preference dataset for image descriptions using unlabeled images. Preferred responses are generated through a step-by-step prompt, while dis-preferred responses are generated from either corrupted images or misleading prompts. To further self-improve reasoning on the extracted visual information, we let the model reuse a small portion of existing instruction-tuning data and append its self-generated image descriptions to the prompts. We validate the effectiveness of STIC across seven different benchmarks, demonstrating substantial performance gains of $4.0\%$ on average while using $70\%$ less supervised fine-tuning data than the current method. Further studies investigate various components of STIC and highlight its potential to leverage vast quantities of unlabeled images for self-training. Code and data are made publicly available.
00footnotetext: Equal contribution.
@@ 1 Introduction
Figure 3: Examples of the self-constructed preference data in STIC.
@@ 2 Related Work
VLMs (Tan and Bansal, 2019; Li et al., 2019, 2020; Kim et al., 2021; Wang et al., 2022b; Bao et al., 2022; Wang et al., 2022a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021a; Shen et al., 2022; Singh et al., 2021), processing both images and text, are pivotal in a wide range of multimodal understanding and reasoning tasks, capable of generating text or encoding multimodal representations. These models have shown increasing proficiency in visual perception and textual reasoning, and are also capable of following complex instructions (OpenAI, 2023b; Team et al., 2023). Recent advancements in the field have been propelled by the availability of open-source large language models (LLMs) (Touvron et al., 2023a, b; Jiang et al., 2023) and innovative image encoders (Radford et al., 2021; Li et al., 2022). For instance, LLaVA (Liu et al., 2023b) combines a vision encoder from CLIP (Radford et al., 2021) with the Vicuna LLM  (Chiang et al., 2023b), and has been further fine-tuned on vision-language instruction-following datasets. The recent development of LVLMs has significantly expanded the scale and diversity of VL instruction-following data, including models such as LLaMA-Adapter-V2 (Gao et al., 2023b), Qwen-VL (Bai et al., 2023), InternVL (Chen et al., 2023b), InstructBLIP (Dai et al., 2024), SPHINX-X (Gao et al., 2024), Claude-3 (Anthropic, 2024), MM1 (McKinzie et al., 2024), and Grok-1.5V (xAI, 2024). In this work, we focus on enhancing the visual perception and mathmatical reasoning capabilities of LVLMs by efficiently aligning them with purely unsupervised data.
Alignment fine-tuning.
@@
Subsequent to supervised fine-tuning (SFT), alignment fine-tuning has emerged as a prominent approach to further enhance the performance of LLMs by aligning them with human preferences (Ouyang et al., 2022; Casper et al., 2023). Early efforts utilized on-policy reinforcement learning (RL) methods, such as proximal policy optimization (PPO) (Schulman et al., 2017), to train a reward model based on preference data (Bai et al., 2022; Touvron et al., 2023a). With the notable introduction of direct policy optimization (DPO) (Rafailov et al., 2023), a new line of research emphasizes direct learning from human preferences without relying on an explicit reward model (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Zheng et al., 2024). Another prominent direction is iterative preference fine-tuning, which has proven effective in enhancing model performance by repeatedly optimizing on newly generated preference pairs in each iteration (Adolphs et al., 2023; Xu et al., 2023; Xiong et al., 2023; Pang et al., 2024). While substantial research has focused on alignment fine-tuning for LLMs, efforts to adapt these techniques for LVLMs have been significantly limited. Initial attempts involve constructing preference datasets using human-labeled data (Sun et al., 2023) or GPT-4 generations for fine-tuning with a DPO loss (Zhou et al., 2024).
Self-improving language models.
High-quality data, including human-crafted and advanced AI generated content, has been demonstrated to significantly enhance the performance of LLMs on various tasks (Josifoski et al., 2023; Taori et al., 2023; Chiang et al., 2023a; Li et al., 2023c). Although, acquiring such high-quality data is often prohibitively expensive. To circumvent the costs associated with obtaining human-annotated or expertly curated data, researchers have shifted their focus to leveraging data generated by the target model itself, exploring ways of self-improvement (Chen et al., 2024; Yuan et al., 2024; Fränken et al., 2024; Rosset et al., 2024). Recent studies have also emphasized the rephrasing capabilities of LLMs, which either enhance their own response quality (Deng et al., 2023; Prasad et al., 2023) or augment synthetic data for self-supervised fine-tuning (Kim et al., 2023). To the best of our knowledge, our work is the first to explore the potential for self-improvement in LVLMs, specifically focusing on the vision modality and emphasizing the self-improvement of image comprehension capabilities.
@@ 3 Problem Setting and Preliminaries
Notation.
We use lower case letters and lower case bold face letters to denote scalars and vectors. We use the symbol $p$ to represent the probability of an LLM’s response. And we denote the sequence of tokens generated from the LLM before the $t$-th token as $\mathbf{y}_{<t}=[y_{1},\ldots,y_{t-1}]$ for $t>1$.
Generative vision language models.
LVLM typically consists of three components: a vision encoder $f(\cdot)$, a projection network $g(\cdot)$, and an LLM $p_{\bm{\theta}}$ parameterized by $\bm{\theta}$. The model processes an image input $\mathbf{e}$ along with a text sequence $\mathbf{x}=[x_{1},\ldots,x_{n}]$ as the prompt to generate a corresponding response $\mathbf{y}=[y_{1},\ldots,y_{m}]$, where $x_{i}$ and $y_{j}$ represent individual tokens from the vocabulary of the LLM. The image is therefore converted into visual tokens within the language token space by the vision encoder and the projection network, producing $\mathbf{v}=[v_{1},\ldots,v_{k}]=f\circ g(\mathbf{e})$. The response $\mathbf{y}$ is then considered as a sample from the conditional probability distribution $p_{\bm{\theta}}(\cdot|\mathbf{v},\mathbf{x})$. As a Markov process, the conditional probability distribution $p_{\bm{\theta}}(\mathbf{y}|\mathbf{v},\mathbf{x})$ can be decomposed as
@@
Alignment fine-tuning.
To improve LLM alignment with human preferences, RL fine-tuning (Bai et al., 2022; Gao et al., 2023a) is typically employed after supervised fine-tuning (SFT). This process involves a reward function $r(\mathbf{x},\mathbf{y})$ for a given sequence pair $(\mathbf{x},\mathbf{y})$. The more preferred response $\mathbf{y}$ is expected to result in a higher reward $r(\mathbf{x},\mathbf{y})$, where the corresponding objective is to maximize the following:
@@
where $\mathbf{x}\sim\mathcal{D}$ is sampled from a given distribution $\mathcal{D}$ and the KL regularization term prevents the new model $p_{\bm{\theta}}$ from deviating too much from the reference model $p_{\mathrm{ref}}$, with $\lambda>0$ as the regularization parameter. Training the reward function is challenging in practice, but direct preference optimization (DPO) (Rafailov et al., 2023) simplifies this process using a predefined preference dataset $S_{\text{pref}}=\big{\{}(\mathbf{x}^{(i)},\mathbf{y}^{(i)}_{w},\mathbf{y}^{(i)}_{l})\big{\}}_{i\in[N]}$, where $\mathbf{y}^{(i)}_{w}$ denotes the preferred response and $\mathbf{y}^{(i)}_{l}$ denotes the dispreferred response given the same prompt $\mathbf{x}^{(i)}$. The objective function is then formulated as
@@
Self-play fine-tuning.
Notably, SPIN (Chen et al., 2024) shares a similar objective formulation while eliminating the need for a preference dataset by considering the model’s own generation as a “dispreferred” response with an iterative self-play mechanism:
$L_{\mathrm{SPIN}}(\bm{\theta},\bm{\theta}_{\mathrm{ref}})=\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim S_{\text{SFT}},\mathbf{y}^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}\bigg{[}\ell\bigg{(}\lambda\log\frac{p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}-\lambda\log\frac{p_{\bm{\theta}}(\mathbf{y}^{\prime}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}^{\prime}|\mathbf{x})}\bigg{)}\bigg{]},$ (3.4)
given a SFT dataset $S_{\mathrm{SFT}}=\{(\mathbf{x},\mathbf{y})\}_{i=1}^{n}$. Here, $p_{\bm{\theta}_{t}}$ represents the LLM in SPIN’s previous iteration.
@@ 4 Our Method: STIC
  Input: Unlabeled image dataset: $\{\mathbf{v}^{(i)}\}_{i\in[N]}$. Image captioning prompt set: $P=\{\mathbf{x}^{(i)}\}_{i\in[M_{1}]}$. Hallucination prompt set: $P_{\mathrm{hallu}}=\{\mathbf{x}_{\mathrm{hallu}}^{(i)}\}_{i\in[M_{2}]}$. Image corruption $h(\cdot)$. Well-curated captioning prompt: $\mathbf{x}_{g}$. LVLM parameterized by $\bm{\theta}_{0}$: $p_{\bm{\theta}_{0}}$.
@@
     Randomly sample a number $n\in(0,1)$.
     Randomly sample $\mathbf{x}\sim\{\mathbf{x}^{(i)}\}_{i\in[M]}$.
     Generate preferred response $\mathbf{y}_{g}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{v}^{(i)},\mathbf{x}_{g})$.
@@
        Generate dispreferred response $\mathbf{y}_{b}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{v}^{(i)},\mathbf{x}_{b})$.
@@
        Generate dispreferred response $\mathbf{y}_{b}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{v}_{b}^{(i)},\mathbf{x}^{(i)})$.
@@
Stage 1: Image comprehension self-training.
The process begins with a self-constructed preference dataset from the base LVLM, which we aim to improve through fine-tuning. The dataset contains paired preference data for image descriptions:
@@
Stage 2: Description-infused fine-tuning.
In the second stage, we further fine-tune the self-trained LVLM to leverage self-generated high-quality image descriptions for instruction-following tasks, and thus help ground its reasoning ability on self-generated descriptions. To achieve this, we randomly select a small subset of data from the model’s instruction fine-tuning dataset already used during SFT. We then infuse the instructions in this subset with model-generated image descriptions as follows:
@@
The original ground-truth completions remain unchanged. We then fine-tune the LVLM for one epoch on this small description-infused subset. This fine-tuning step ensures that the model effectively integrates visual information into its responses, thereby enhancing its ability to handle a variety of vision-language reasoning tasks. During inference, optionally, we can let the model self-augment its prompt for downstream vision-language reasoning tasks by describing the image before answering the question.
@@
     Add $\big{(}[\mathbf{x}_{\text{des}},\mathbf{x}^{(i)}],\mathbf{y}^{(i)}\big{)}$ to $D_{\text{des}}$.
@@ 5 Experiments
In this section, we present the experiment results of STIC across seven visual question answering (VQA) benchmarks. We demonstrate that STIC effectively and substantially improves LVLM’s performance across different VQA tasks using a self-constructed preference dataset without external labels.
@@ 5.1 Experiment Setup
Evaluation.
We consider the widely used benchmarks for LVLM evaluation across different domains including: ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), ChartQA (Masry et al., 2022), LLaVA-Bench (Liu et al., 2023a), MMBench (Liu et al., 2023c), MM-Vet (Yu et al., 2023), and MathVista (Lu et al., 2024). Specifically, ScienceQA focuses on scientific question answering and MathVista focuses on math reasoning with visual information. TextVQA consists of images with text-rich contents and ChartQA with visual charts. Lastly, LLaVA-Bench, MMBench, and MM-Vet are three recent benchmarks to comprehensively evaluate a model’s capabilities in a wide range of tasks and evaluation criteria. We use the evaluation scripts provided by LLaVA (Liu et al., 2023a) to obtain the results for both our base model and after using STIC to ensure a fair comparison.

### v2

@@ Abstract
Large vision language models (LVLMs) integrate large language models (LLMs) with pre-trained vision encoders, thereby activating the model’s perception capability to understand image inputs and conduct subsequent reasoning for different queries. Improving this capability requires high-quality vision-language data, which is costly and labor-intensive to acquire. Self-training approaches have been effective in single-modal settings to alleviate the need for labeled data by leveraging model’s own generation. However, effective self-training remains a challenge regarding the unique visual perception and reasoning capability of LVLMs. To address this, we introduce Self-Training on Image Comprehension (STIC), which emphasizes a self-training approach specifically for image comprehension. First, the model self-constructs a preference dataset for image descriptions using unlabeled images. Preferred responses are generated through a step-by-step prompt, while dis-preferred responses are generated from either corrupted images or misleading prompts. To further self-improve reasoning on the extracted visual information, we let the model reuse a small portion of existing instruction-tuning data and append its self-generated image descriptions to the prompts. We validate the effectiveness of STIC across seven different benchmarks, demonstrating substantial performance gains of $4.0\%$ on average while using $70\%$ less supervised fine-tuning data than the current method. Further studies investigate various components of STIC and highlight its potential to leverage vast quantities of unlabeled images for self-training. Code and data are made publicly available on GitHub.
00footnotetext: Equal contribution.00footnotetext: Contribution statement is provided after Section 7.
@@ 1 Introduction
Figure 3: Examples of the self-constructed preference data in STIC.
@@ 2 Related Work
Vision language models (VLMs). VLMs (Tan and Bansal, 2019; Li et al., 2019, 2020; Kim et al., 2021; Wang et al., 2022b; Bao et al., 2022; Wang et al., 2022a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021a; Shen et al., 2022; Singh et al., 2021), processing both images and text, are pivotal in a wide range of multimodal understanding and reasoning tasks, capable of generating text or encoding multimodal representations. These models have shown increasing proficiency in visual perception and textual reasoning, and are also capable of following complex instructions (OpenAI, 2023b; Team et al., 2023). Recent advancements in the field have been propelled by the availability of open-source large language models (LLMs) (Touvron et al., 2023a, b; Jiang et al., 2023) and innovative image encoders (Radford et al., 2021; Li et al., 2022). For instance, LLaVA (Liu et al., 2023b) combines a vision encoder from CLIP (Radford et al., 2021) with the Vicuna LLM  (Chiang et al., 2023b), and has been further fine-tuned on vision-language instruction-following datasets. The recent development of LVLMs has significantly expanded the scale and diversity of VL instruction-following data, including models such as LLaMA-Adapter-V2 (Gao et al., 2023b), Qwen-VL (Bai et al., 2023), InternVL (Chen et al., 2023b), InstructBLIP (Dai et al., 2024), SPHINX-X (Gao et al., 2024), Claude-3 (Anthropic, 2024), MM1 (McKinzie et al., 2024), and Grok-1.5V (xAI, 2024). In this work, we focus on enhancing the visual perception and mathmatical reasoning capabilities of LVLMs by efficiently aligning them with purely unsupervised data.
@@
Alignment fine-tuning. Subsequent to supervised fine-tuning (SFT), alignment fine-tuning has emerged as a prominent approach to further enhance the performance of LLMs by aligning them with human preferences (Ouyang et al., 2022; Casper et al., 2023). Early efforts utilized on-policy reinforcement learning (RL) methods, such as proximal policy optimization (PPO) (Schulman et al., 2017), to train a reward model based on preference data (Bai et al., 2022; Touvron et al., 2023a). With the notable introduction of direct policy optimization (DPO) (Rafailov et al., 2023), a new line of research emphasizes direct learning from human preferences without relying on an explicit reward model (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Zheng et al., 2024). Based on DPO, a line of research focus on iterative fine-tuning that repeatedly optimizes on newly generated preference pairs in each iteration (Adolphs et al., 2023; Xu et al., 2023; Xiong et al., 2023; Pang et al., 2024). While substantial research has focused on alignment fine-tuning for LLMs, efforts to adapt these techniques for LVLMs have been significantly limited. Initial attempts involve constructing preference datasets using human-labeled data (Sun et al., 2023) or GPT-4 generations for fine-tuning with a DPO loss (Zhou et al., 2024). Concurrent works (Pi et al., 2024; Zhou et al., 2024) begin to focus on generating preference dataset of LVLMs, while our method distinguishes itself with the unique preference prompt set.
Self-training. Traditional self-supervised training schemes (He et al., 2019; Xie et al., 2020; Wei et al., 2020; Zoph et al., 2020; Sohn et al., 2020; Ghiasi et al., 2021; Kang et al., 2023) leverage trained models to generate labels for unlabeled data and incorporate these self-labeled examples into training as a form of data augmentation. While both classical self-training schemes and our approach share the fundamental goal of effectively utilizing unlabeled data to enhance model performance, our method differs in its focus on vision LLMs, maintaining an LLM as the backbone architecture. Rather than optimizing image representations, our approach aims to generate synthetic data that enables the LLM to produce higher-quality responses to image queries.
@@ 3 Problem Setting and Preliminaries
Notation. We use lower case letters and lower case bold face letters to denote scalars and vectors. We use the symbol $p$ to represent the probability of an LLM’s response. And we denote the sequence of tokens generated from the LLM before the $t$-th token as $\mathbf{y}_{<t}=[y_{1},\ldots,y_{t-1}]$ for $t>1$.
Generative vision language models. LVLM typically consists of three components: a vision encoder $f(\cdot)$, a projection network $g(\cdot)$, and an LLM $p_{\bm{\theta}}$ parameterized by $\bm{\theta}$. The model processes an image input $\mathbf{e}$ along with a text sequence $\mathbf{x}=[x_{1},\ldots,x_{n}]$ as the prompt to generate a corresponding response $\mathbf{y}=[y_{1},\ldots,y_{m}]$, where $x_{i}$ and $y_{j}$ represent individual tokens from the vocabulary of the LLM. The image is therefore converted into visual tokens within the language token space by the vision encoder and the projection network, producing $\mathbf{v}=[v_{1},\ldots,v_{k}]=f\circ g(\mathbf{e})$. The response $\mathbf{y}$ is then considered as a sample from the conditional probability distribution $p_{\bm{\theta}}(\cdot|\mathbf{v},\mathbf{x})$. As a Markov process, the conditional probability distribution $p_{\bm{\theta}}(\mathbf{y}|\mathbf{v},\mathbf{x})$ can be decomposed as
@@
Alignment fine-tuning. To improve LLM alignment with human preferences, RLHF (Bai et al., 2022; Gao et al., 2023a) is typically employed after supervised fine-tuning (SFT). This process involves training a reward function $r(\mathbf{x},\mathbf{y})$ for a given sequence pair $(\mathbf{x},\mathbf{y})$, based on a preference dataset $S_{\text{pref}}=\big{\{}(\mathbf{x}^{(i)},\mathbf{y}^{(i)}_{w},\mathbf{y}^{(i)}_{l})\big{\}}_{i\in[N]}$, where $\mathbf{y}^{(i)}_{w}$ denotes the preferred response and $\mathbf{y}^{(i)}_{l}$ denotes the dispreferred response for the same prompt $\mathbf{x}^{(i)}$. The more preferred response $\mathbf{y}$ is expected to yield a higher reward $r(\mathbf{x},\mathbf{y})$. The optimization objective for RL fine-tuning is formulated as follows:
@@
where $\mathbf{x}\sim\mathcal{D}$ is sampled from a given distribution $\mathcal{D}$ and the KL regularization term prevents the new model $p_{\bm{\theta}}$ from deviating too much from the reference model $p_{\mathrm{ref}}$, with $\lambda>0$ as the regularization parameter. This objective is typically optimized using online RL algorithms (Schulman et al., 2017; Ahmadian et al., 2024). However, these algorithms are often computationally intensive, requiring reward model training and sampling from the policy model in training. Direct preference optimization (DPO) (Rafailov et al., 2023) simplifies this process by defining an implicit reward model with the policy and reference model. Instead of training a separate reward model, DPO directly optimizes the policy model on the preference dataset. The corresponding objective function is:
@@ 4 Our Method: STIC
  Input: Unlabeled image dataset: $\{\mathbf{v}^{(i)}\}_{i\in[N]}$. Image captioning prompt set: $P=\{\mathbf{x}^{(i)}\}_{i\in[M_{1}]}$, where $M_{1}$ is the size of the set. Hallucination prompt set: $P_{\mathrm{hallu}}=\{\mathbf{x}_{\mathrm{hallu}}^{(i)}\}_{i\in[M_{2}]}$, where $M_{2}$ is the size of the set. Image corruption $h(\cdot)$. Well-curated captioning prompt: $\mathbf{x}_{g}$. LVLM parameterized by $\bm{\theta}_{0}$: $p_{\bm{\theta}_{0}}$.
@@
     Randomly sample a number $n\in[0,1]$.
     Randomly sample $\mathbf{x}\sim P$.
     Generate preferred response $\mathbf{y}_{g}\sim p_{\bm{\theta}_{0}}(\cdot|\mathbf{v}^{(i)},\mathbf{x}_{g})$.
@@
        Generate dispreferred response $\mathbf{y}_{b}\sim p_{\bm{\theta}_{0}}(\cdot|\mathbf{v}^{(i)},\mathbf{x}_{b})$.
@@
        Generate dispreferred response $\mathbf{y}_{b}\sim p_{\bm{\theta}_{0}}(\cdot|\mathbf{v}_{b}^{(i)},\mathbf{x})$.
@@
Stage 1: Image comprehension self-training. The process begins with a self-constructed preference dataset from the base LVLM, which we aim to improve through fine-tuning. The dataset contains paired preference data for image descriptions:
@@
Prompt design. Our prompt design for the well-crafted prompt focuses on quality and diversity. We use GPT-4 to generate and sample multiple initial prompts, which are then refined through human filtering. To ensure effectiveness, we test these prompts on MSCOCO samples, verifying their ability to produce well-structured and relevant responses from the model. The bad prompts are sampled from GPT-4 and, in contrast, designed to elicit inaccurate descriptions by setting up a slightly different task (describe objects that would logically exist in the image) for the model. We thus work under the assumption that responses generated from prompts that have differences in human preference lead to responses of the same preference with high probability. The key is that the discrepancy between good and bad prompts should result in pairs of responses that share the same implicit preference with high probability, which is sufficient for effective DPO training.
Stage 2: Description-infused fine-tuning. In the second stage, we further fine-tune the self-trained LVLM to leverage self-generated high-quality image descriptions for instruction-following tasks, and thus help ground its reasoning ability on self-generated descriptions. To achieve this, we randomly select a small subset of data from the model’s instruction fine-tuning dataset already used during SFT. We then infuse the instructions in this subset with model-generated image descriptions as follows:
@@
The original ground-truth completions remain unchanged. We then fine-tune the LVLM for one epoch on this small description-infused subset. This fine-tuning step ensures that the model effectively integrates visual information into its responses, thereby enhancing its ability to handle a variety of vision-language reasoning tasks.
Describe and Respond. During inference, optionally, we can let the model self-augment its prompt for downstream vision-language reasoning tasks by describing the image before answering. Rather than generating an immediate response, we first elicit an image description, which is then concatenated with the original question to produce a more informed answer.
@@
     Add $\big{(}[\mathbf{y}_{\text{des}},\mathbf{x}^{(i)}],\mathbf{y}^{(i)}\big{)}$ to $D_{\text{des}}$.
@@ 5 Experiments
In this section, we present the experiment results of STIC across seven visual question answering (VQA) benchmarks. We demonstrate that STIC effectively and substantially improves LVLM’s performance across different VQA tasks using a self-constructed preference dataset without labels.
@@ 5.1 Experiment Setup
Evaluation. We consider the widely used benchmarks for LVLM evaluation across different domains including: ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), ChartQA (Masry et al., 2022), LLaVA-Bench (Liu et al., 2023a), MMBench (Liu et al., 2023c), MM-Vet (Yu et al., 2023), and MathVista (Lu et al., 2024). Specifically, ScienceQA focuses on scientific question answering and MathVista focuses on math reasoning with visual information. TextVQA consists of images with text-rich contents and ChartQA with visual charts. Lastly, LLaVA-Bench, MMBench, and MM-Vet are three recent benchmarks to comprehensively evaluate a model’s capabilities in a wide range of tasks and evaluation criteria. We use the evaluation scripts provided by LLaVA (Liu et al., 2023a) to obtain the results for both our base model and after using STIC to ensure a fair comparison.